OpenStudy (bubbles5705):

How to remember maths. So that when you get a test you can at least get a 60%.

OpenStudy (anonymous):

the best way is to practice and practice ,,above all before the test

OpenStudy (anonymous):

Exactly as ^ said. What level of math are you up to? Some levels require a lot more memorization than others.

OpenStudy (bubbles5705):

Ok Thanks

OpenStudy (anonymous):

remember jokes that has to do with the rules on how to do things or defientions

OpenStudy (anonymous):

Yeah. You might remember seeing one about \(\sin^2 \theta + \cos^2 \theta = 1\)

OpenStudy (bubbles5705):

Yep